### v3

@@ Abstract
Protein structure is key to understanding protein function and is essential for progress in bioengineering, drug discovery, and molecular biology. Recently, with the incorporation of generative AI, the power and accuracy of computational protein structure prediction/design have been improved significantly. However, ethical concerns such as copyright protection and harmful content generation (biosecurity) pose challenges to the wide implementation of protein generative models. Here, we investigate whether it is possible to embed watermarks into protein generative models and their outputs for copyright authentication and the tracking of generated structures. As a proof of concept, we propose a two-stage method FoldMark as a generalized watermarking strategy for protein generative models. FoldMark first pretrain watermark encoder and decoder, which can minorly adjust protein structures to embed user-specific information and faithfully recover the information from the encoded structure. In the second step, protein generative models are fine-tuned with watermark Low-Rank Adaptation (LoRA) modules to preserve generation quality while learning to generate watermarked structures with high recovery rates. Extensive experiments are conducted on open-source protein structure prediction models (e.g., ESMFold and MultiFlow) and de novo structure design models (e.g., FrameDiff and FoldFlow) and we demonstrate that our method is effective across all these generative models. Meanwhile, our watermarking framework only exerts a negligible impact on the original protein structure quality and is robust under potential post-processing and adaptive attacks.
@@ Introduction
Similar problems have occurred in text and image generation. For example, Large language models (LLMs) such as ChatGPT can be used to create fake news and to cheat on academic writing [16, 17, 18]. The latest text-to-image models such as Stable Diffusion [19] and DALL·E 3 [20] enable users to create photo-realistic images like deep fakes [21] for illegal purposes. As a result, there is growing consensus that the ability to detect, track, and audit the use of AI-generated content is essential for harm reduction and regulation [17, 22]. Recently, the watermark becomes one of the most promising protection strategy, which embeds hidden patterns in the generated content and is imperceptible to humans, while making the embedded information algorithmically identifiable. Although watermark has been applied for LLM [23, 24, 25, 26, 27] and text-to-image models [28, 29, 30, 31], extending these methods to protein structure data presents unique challenges. Unlike text and images, protein structures are highly sensitive to minute changes [2, 32, 33], and embedding watermarks without disrupting the biological functionality or stability of the proteins is a complex task. Moreover, protein structures exhibit complex geometrical symmetries, making traditional watermarking methods less effective due to the requirement for equivariance [4, 34].
In this paper, as a proof of concept, we propose FoldMark, a generalized watermarking method for protein generative models, e.g., AlphaFold2 [2] and RFDiffusion [4]. FoldMark builds upon pretrained protein generative models and generally has two training stages. In the first stage, SE(3)-equivariant watermark encoder and decoder are pretrained to learn how to embed watermark information into protein structure without compromising the structural quality. Specifically, the encoder takes the watermark code and learns to construct the watermark-conditioned structure; the decoder takes the watermarked structure and predicts the embedded watermark. In the second stage, we propose watermark LoRA [35], which flexibly encodes the given watermark code and merges it into the original model weights, without changing or adding extra model architecture. The protein generative model is fine-tuned with two objectives, named message retrieval loss and consistency loss. The message retrieval loss ensures the effective embedding of watermarks into the generated structure, allowing for successful retrieval of the embedded watermark codes. Meanwhile, the consistency loss ensures that the inclusion of the watermark has minimal impact on the overall quality of the protein structure.
@@ Methods
Figures 2 and 3 provide an overview of our method, FoldMark, which consists of two main stages: Watermark Encoder/Decoder Pretraining and Prior-Preserving Finetuning. The pretraining stage enables the watermark encoder and decoder to learn how to embed watermark information into the structure space and accurately extract it. The finetuning stage equips pretrained protein generative models with watermarking capabilities while preserving their original generative performance (prior-preserving). FoldMark is a versatile method that can be applied to various mainstream protein structure generative models. We use a diffusion-based model as an example and present the details of FoldMark below.
@@
Prior-preserving Finetuning
Instead of finetuning all the parameters of the generative model, we selectively fine-tune part of the protein generative model with watermark LoRA and the watermark decoder as shown in Figure 3. The other parameters including the watermark embedder $\mathcal{P}$ and the reference model are kept unchanged. We discuss the details of watermark module in the next subsection.
Here we take the diffusion-based protein generative model (e.g., FrameDiff [37] and RFDiffusion [4]) as an example to construct the fine-tuning loss. The diffusion model typically involves two critical components known as the forward and backward process, where the forward process gradually noises the original protein structure $\mathbf{x}_{0}$ into $\mathbf{x}_{t}$ for $t\in\{1,\cdots,T\}$ and the model learns to predict the original structure $\epsilon_{\theta}(\mathbf{x}_{t})$ based on $\mathbf{x}_{t}$. There are two losses in the fine-tuning: the consistency loss for regularization and the message retrieval loss to encourage correct watermark retrieval. In the consistency loss $\mathcal{L}_{c}$, the prediction of the fine-tuned model is compared with the original pretrained model so that the finetuned model weights will not deviate too much from the original ones. For the watermark retrieval loss $\mathcal{L}_{m}$, we take a single reverse step with respect to $\mathbf{x}_{t}$ to obtain $\hat{\mathbf{x}}_{t}=(\mathbf{x}_{t}-\sqrt{1-\bar{\alpha_{t}}}\epsilon_{\theta}(\mathbf{x}_{t}))/\sqrt{\bar{\alpha_{t}}}$, and then feed it into the decoder to predict the watermark code. In sum, we incorporate both optimization objectives above and formulate the prior-preserving finetuning loss as:
@@
Watermark LoRA
To save the computation costs of fine-tuning and flexibly embed watermark information in the generation process, we proposed the Watermark LoRA strategy. The computation formula for LoRA can be expressed as $\Delta\mathbf{W}=\mathbf{A}\times\mathbf{E}\times\mathbf{B}$, where $\mathbf{A}\in\mathbb{R}^{n\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times m}$, and $\mathbf{E}\in\mathbb{R}^{r\times r}$ serves as the embedding matrix. In this context, $r$ represents the rank of the matrix, satisfying $r\ll n,m$. When $\mathbf{E}$ is set as an identity matrix, the computation aligns with the vanilla LoRA setup [35].
To input the watermark information into the fine-tuned model, We utilize a mapping function that converts a watermark code of length $l$ into a vector of length $r$. For each $i$-th bit of the code, an embedding vector $\mathbf{I}_{i}$ represents the binary state 1, while $\mathbf{0}$ represents the binary state 0, where $\mathbf{I}_{i},\mathbf{0}\in\mathbb{R}^{r}$. This mapping function, $h_{i}:\{0,1\}\rightarrow\mathbb{R}^{r}$, is defined as
$h_{i}(b_{i})=\begin{cases}\mathbf{I}_{i},&\text{if }b_{i}=1,\\ \mathbf{0},&\text{otherwise}.\end{cases}$ (3)
For a watermark code $\textbf{m}=\{b_{0},b_{1},\ldots,b_{l}\}$, the embedding matrix $\mathbf{E}$ is constructed as
$\mathbf{S}=\mathrm{diag}\left(\textbf{1}+\frac{1}{\sqrt{l}}\sum_{i=1}^{l}h_{i}(b_{i})\right).$ (4)
During the generation process, when adding a watermark to the model for a given watermark code, we pass the watermark through the mapper as mentioned earlier to obtain the embedding matrix $\mathbf{E}$. We then calculate the added model weights $\Delta\mathbf{W}$ and incorporate it into the original model weights using the formula: $\mathbf{W}_{\text{watermarked}}=\mathbf{W}+\alpha\Delta\mathbf{W}$.
$\mathbf{I}_{i}$ can be initialized from a standard normal distribution. In FoldMark, we propose using orthogonal initialization to maximize the differences between vectors. Specifically, for any $\mathbf{I}_{j}$ and $\mathbf{I}_{k}$, where $j,k\in[0,l]$, we ensure that $\mathbf{I}_{j}\cdot\mathbf{I}_{k}=\mathbf{0}$. During model training, we fixed the mapper’s weights, as fine-tuning them showed no significant improvement.
@@
Our FoldMark model is pretrained for 20 epochs and fine-tuned for 10 epochs with Adam [43] optimizer, where the learning rate is 0.0001, and the max batch size is 64. We use the batching strategy from FrameDiff [44] of combining proteins with the same length into the same batch to remove extraneous padding. In the watermark LoRA, the rank is set as 16 in the default setting. $\gamma$ and $\eta$ are set as 2. We report the results corresponding to the checkpoint with the best validation loss. It takes less than 48 hours to finish the whole training process on 1 Tesla A100 GPU. More hyperparameter settings are listed in Table. LABEL:tab:hyperparam.
@@
Ablation studies of FoldMark are here.
Watermark Encoder-Decoder architecture

### v4

@@ Abstract
Protein structure is key to understanding protein function and is essential for progress in bioengineering, drug discovery, and molecular biology. Recently, with the incorporation of generative AI, the power and accuracy of computational protein structure prediction/design have been improved significantly. However, ethical concerns such as copyright protection and harmful content generation (biosecurity) pose challenges to the wide implementation of protein generative models. Here, we investigate whether it is possible to embed watermarks into protein generative models and their outputs for copyright authentication and the tracking of generated structures. As a proof of concept, we propose a two-stage method FoldMark as a generalized watermarking strategy for protein generative models. FoldMark first pretrain watermark encoder and decoder, which can minorly adjust protein structures to embed user-specific information and faithfully recover the information from the encoded structure. In the second step, protein generative models are fine-tuned with Watermark-conditioned Low-Rank Adaptation (WaterLoRA) modules to preserve generation quality while learning to generate watermarked structures with high recovery rates. Extensive experiments are conducted on open-source protein structure prediction models (e.g., ESMFold and MultiFlow) and de novo structure design models (e.g., FrameDiff and FoldFlow) and we demonstrate that our method is effective across all these generative models. Meanwhile, our watermarking framework only exerts a negligible impact on the original protein structure quality and is robust under potential post-processing and adaptive attacks.
@@ Introduction
Similar problems have occurred in text and image generation. For example, Large language models (LLMs) such as ChatGPT can be used to create fake news and to cheat on academic writing [16, 17, 18]. The latest text-to-image models such as Stable Diffusion [19] and DALL·E 3 [20] enable users to create photo-realistic images like deep fakes [21] for illegal purposes. As a result, there is growing consensus that the ability to detect, track, and audit the use of AI-generated content is essential for harm reduction and regulation [17, 22]. Recently, the watermark becomes one of the most promising protection strategy, which embeds hidden patterns in the generated content and is imperceptible to humans, while making the embedded information algorithmically identifiable. Although watermark has been applied for LLM [23, 24, 25, 26, 27] (e.g., SynthID-Text [27]) and text-to-image models [28, 29, 30, 31] (e.g., AquaLoRA [30]), extending these methods to protein structure data presents unique challenges. Unlike text and images, protein structures are highly sensitive to minute changes [2, 32, 33], and embedding watermarks without disrupting the biological functionality or stability of the proteins is a complex task. Moreover, protein structures exhibit complex geometrical symmetries, making traditional watermarking methods less effective due to the requirement for equivariance [4, 34].
In this paper, as a proof of concept, we propose FoldMark, a generalized watermarking method for protein generative models, e.g., AlphaFold2 [2] and RFDiffusion [4]. FoldMark builds upon pretrained protein generative models and generally has two training stages. In the first stage, SE(3)-equivariant watermark encoder and decoder are pretrained to learn how to embed watermark information into protein structure without compromising the structural quality. Specifically, the encoder takes the watermark code and learns to construct the watermark-conditioned structure; the decoder takes the watermarked structure and predicts the embedded watermark. In the second stage, we propose WaterLoRA [35], which flexibly encodes the given watermark code and merges it into the original model weights, without changing or adding extra model architecture. The protein generative model is fine-tuned with two objectives, named message retrieval loss and consistency loss. The message retrieval loss ensures the effective embedding of watermarks into the generated structure, allowing for successful retrieval of the embedded watermark codes. Meanwhile, the consistency loss ensures that the inclusion of the watermark has minimal impact on the overall quality of the protein structure.
@@ Methods
Figures 2 and 3 provide an overview of our method. Inspired by previous works [28, 30, 29], FoldMark consists of two main stages: Watermark Encoder/Decoder Pretraining and Consistency-Preserving Finetuning. The pretraining stage enables the watermark encoder and decoder to learn how to embed watermark information into the structure space and accurately extract it. The finetuning stage equips pretrained protein generative models with watermarking capabilities while preserving their original generative performance (Consistency-preserving). FoldMark is a versatile method that can be applied to various mainstream protein structure generative models. We use a diffusion-based model as an example and present the details of FoldMark below.
@@
Consistency-preserving Finetuning
Instead of finetuning all the parameters of the generative model, we selectively fine-tune part of the protein generative model with WaterLoRA and the watermark decoder as shown in Figure 3. The other parameters including the watermark embedder $\mathcal{P}$ and the reference model are kept unchanged. We discuss the details of watermark module in the next subsection.
Here we take the diffusion-based protein generative model (e.g., FrameDiff [37] and RFDiffusion [4]) as an example to construct the fine-tuning loss. The diffusion model typically involves two critical components known as the forward and backward process, where the forward process gradually noises the original protein structure $\mathbf{x}_{0}$ into $\mathbf{x}_{t}$ for $t\in\{1,\cdots,T\}$ and the model learns to predict the original structure $\epsilon_{\theta}(\mathbf{x}_{t})$ based on $\mathbf{x}_{t}$. There are two losses in the fine-tuning: the consistency loss for regularization and the message retrieval loss to encourage correct watermark retrieval. In the consistency loss $\mathcal{L}_{c}$, the prediction of the fine-tuned model is compared with the original pretrained model so that the finetuned model weights will not deviate too much from the original ones. For the watermark retrieval loss $\mathcal{L}_{m}$, we take a single reverse step with respect to $\mathbf{x}_{t}$ to obtain $\hat{\mathbf{x}}_{t}=(\mathbf{x}_{t}-\sqrt{1-\bar{\alpha_{t}}}\epsilon_{\theta}(\mathbf{x}_{t}))/\sqrt{\bar{\alpha_{t}}}$, and then feed it into the decoder to predict the watermark code. In sum, we incorporate both optimization objectives above and formulate the consistency-preserving finetuning loss as:
@@
Watermark-conditioned LoRA (WaterLora)
To save the computation costs of fine-tuning and flexibly embed watermark information in the generation process, we propose the Watermark-conditioned LoRA (WaterLoRA) strategy, inspired by previous works [35, 41, 42, 30, 43] on parameter-efficient fine-tuning and modulation. The computation formula for WaterLoRA can be expressed as:
$\Delta\mathbf{W}(\mathbf{m})=\mathbf{G}(\mathbf{m})\odot(\mathbf{A}\times\mathbf{B}),$
where $\mathbf{A}\in\mathbb{R}^{n\times r}$ and $\mathbf{B}\in\mathbb{R}^{r\times m}$ are the low-rank matrices, and $\mathbf{G}\in\mathbb{R}^{n}$ is the gating vector derived from the watermark code. The operator $\odot$ denotes element-wise multiplication, where $\mathbf{G}$ modulates the rows of $\mathbf{A}\times\mathbf{B}$. This formulation maintains efficiency while allowing flexible incorporation of watermark information.
To input the watermark information into the fine-tuned model, we utilize an adapter layer that converts a watermark code of length $l$ into a gating vector $\mathbf{G}$. Specifically, the watermark code $\mathbf{m}=\{b_{0},b_{1},\ldots,b_{l}\}$ is passed through a linear transformation defined as:
$\mathbf{G}(\mathbf{m})=\mathbf{W}_{g}\cdot\mathbf{m}+\mathbf{b}_{g},$
where $\mathbf{W}_{g}\in\mathbb{R}^{n\times l}$ is the linear transformation matrix, and $\mathbf{b}_{g}\in\mathbb{R}^{n}$ is the bias vector. Here, $b_{i}\in\{0,1\}$ represents the binary state of the $i$-th bit in the watermark code. The gating vector $\mathbf{G}$ modulates the LoRA weight updates by scaling the rows of the low-rank update $\mathbf{A}\times\mathbf{B}$.
During the generation process, when embedding a watermark into the model, we compute the gating vector $\mathbf{G}$ based on the watermark code. The resulting LoRA weight update $\Delta\mathbf{W}$ is added to the original model weights to produce the watermarked model weights:
$\mathbf{W}_{\text{watermarked}}=\mathbf{W}+\alpha\Delta\mathbf{W}(\mathbf{m}),$
where $\alpha$ is a scaling factor controlling the impact of the watermark on the model weights. FoldMark apply WaterLora to all linear and attention layers for effective watermarking.
@@
Our FoldMark model is pretrained for 20 epochs and fine-tuned for 10 epochs with Adam [46] optimizer, where the learning rate is 0.0001, and the max batch size is 64. We use the batching strategy from FrameDiff [47] of combining proteins with the same length into the same batch to remove extraneous padding. In the WaterLoRA, the rank is set as 16 in the default setting. $\gamma$ and $\eta$ are set as 2. We report the results corresponding to the checkpoint with the best validation loss. It takes less than 48 hours to finish the whole training process on 1 Tesla A100 GPU. More hyperparameter settings are listed in Table. LABEL:tab:hyperparam.
@@
0.1 Comparison with other watermark methods
Traditional watermarking techniques developed for Large Language Models (LLMs) and diffusion models are not directly applicable to protein structure data due to the unique characteristics of protein structures. Similar methods such as WaDiff [29] and AquaLoRA [30], which embed watermarks into the U-Net backbone of Stable Diffusion models for image generation, are effective in the image domain but typically fail in protein generative models. FoldMark addresses this gap by employing optimized pretraining and fine-tuning losses, a customized encoder-decoder architecture, and enhanced watermark-conditioned Low-Rank Adaptation (LoRA) methods, achieving superior performance in protecting protein generative models.
0.2 Ablation studies of FoldMark
0.3 Watermark Encoder-Decoder architecture